2 blocks are connected over a massless pulley. The mass of block A is 10kg and the coefficient of kinetic friction is 0.2. Block A slides down the incline at constant speed. The incline makes an angle of 30° with the base. What is the mass of block B? (Block B isn`t in contact with the wedge)

Grade:11

1 Answers

devki
6 years ago
T-Ff -W
Ff =muek N
=muek(mgcostheta)
W=mgsintheta 
T =mg(sin30-muekcos30)
T-mg=0
m=m(sin30-muekcos30)
m=3.3kg